OpenStudy (anonymous):

Evaluate the integral using substitution method. ∫sec(x/3)dx

OpenStudy (anonymous):

hint: the integral of sec(u) is = ln(sec(u) + tan(u)) so now you can guess what U should be :)

OpenStudy (anonymous):

oh ok thank you ! =)

OpenStudy (anonymous):

here if you want to see all the work: u = x/3 du = dx = integral [ sec(u) du ] this part is fun, my mind was blown when I found out you could do this. It's almost like magic: = integral [ sec(u) * [sec(u) + tan(u) / sec(u) + tan(u)] du] = integral [ sec^2(u) + sec(u)tan(u) / sec(u) + tan(u) du] now lets use a W substitution w = sec(u) + tan(u) dw = sec(u)tan(u) + sec^2(u) du you see how we have a sec^2(u) + sec(u)tan(u) at the top already? = integral [ (1/w) dw] = ln(w) = ln( sec(u) + tan(u)) = ln(sec(x/3) + tan(x/3)) tada!

OpenStudy (anonymous):

oh sorry there should be a 1/3 in the front

OpenStudy (anonymous):

yes i saw that. thank you very much again